OpenStudy (anonymous):

Convert the recurring decimal 6.838383 into an fraction???

OpenStudy (anonymous):

\[ \frac{677}{99} \]

OpenStudy (anonymous):

Exercise for you, convert \[ 6.7845838383838383\cdots \] to rational.

myininaya (myininaya):

\[\text{ let } x=6.838383838...\] \[\text{ then } 100x=683.838838383...\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=6.7845838383838383... DAMN I'm good.

myininaya (myininaya):

now do 100x-x

myininaya (myininaya):

\[99x=683-6\]

myininaya (myininaya):

solve for x

myininaya (myininaya):

\[99x=677 => x=\frac{677}{99}\]

OpenStudy (anonymous):

myin, do you know the one liner trick? ;)

myininaya (myininaya):

@strawberry how do you feel about that?

OpenStudy (anonymous):

pretty confident now

myininaya (myininaya):

i multiply x by 100 because i seen that the 83 was repeating